A unit of time sometimes used in microscopic physics is the shake. One shake is 10 seconds. Compare the number of shakes in a second with the number of seconds in a year, which is more?
1
Expert's answer
2021-10-01T16:56:07-0400

If one shake is 10 seconds, than there are 0.1 shake in one second.

There are


"365\times 24\times 60\times 60 =31536000"

seconds in a year.

Thus, the number of seconds in a year is much greater than the number of shakes in second.